OpenStudy (anonymous):

A family of 5 comes to the amusement park. There are two adults and three children (under the age of 12). They qualify for a special family rate of $174.45. The clerk tells them that a child's ticket always costs $11 less than an adult's ticket. How much is each adult ticket and how much is each child's ticket?

OpenStudy (anonymous):

Let cost of one adult ticket be X $. => cost of one child's ticket = X - 11 $ Now there are 2 adults and 3 children. Therefore 2 X + 3* (X-11) = 174.45 Solve that and you have your answer!

OpenStudy (anonymous):

2x + 3(x-11) = 174.45 where x is cost of adult ticket

OpenStudy (anonymous):

Solve the following for a and c:\[\{2a +3c=174.45,c\text{ = }a-11\}\]\[\{a=\text{$\$$41}.49 \text{ each},c=\text{$\$$30}.49 \text{ each}\} \]